OpenStudy (anonymous):

Linda bought an 8 foot long piece of framing to frame her 2 college awards. One award is 8.5 inches by 11 inches and the other award is 11 inches by 17 inches. Will she have enough framing or will she be short? By how much?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

here the area of the first frame, is: 8.5*11=...? inches^2

OpenStudy (anonymous):

93.5

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

whereas the area of the second frame is: 11*17=...? inches^2

OpenStudy (anonymous):

187

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

and their sum is: 187+93.5=...? inches^2

OpenStudy (anonymous):

2805

OpenStudy (michele_laino):

280.5 inches^2

OpenStudy (michele_laino):

is it ok?

OpenStudy (anonymous):

I think the answer is 1 inch left over

OpenStudy (michele_laino):

8 feet = 12*8 = 96 inches

OpenStudy (michele_laino):

oops.. we have to compute the perimeters of each frame, sorry: perimeter of first frame is: 2*(8.5+11)=39 inches perimeter of second frame is: 2*(11+17)=56 inches. total perimeter = 95 inches. so we have enough framing since Linda bought 96 inches of framing

OpenStudy (anonymous):

okay!:)

OpenStudy (michele_laino):

Thank you! :)